# ChainMPQ: Interleaved Text-Image Reasoning Chains for Mitigating Relation Hallucinations

### v1

@@ 5 CONCLUSION
We propose ChainMPQ (Multi-Perspective Questions guided Interleaved Chain of Image and Text), a training-free framework that enhances the ability of LVLMs to recognize relationships by utilizing accumulated textual and visual memories. ChainMPQ first strengthens the relevant visual tokens using keywords extracted from the question. It then constructs a set of complementary questions from different perspectives, each closely centered on the subject, the object, and their relationship. These questions are sequentially fed into the model, with both textual and visual memories from earlier rounds propagated to subsequent ones, thereby forming an interleaved chain of image and text. Experimental results demonstrate that ChainMPQ effectively reduces relation hallucinations and improves factuality across multiple LVLMs and benchmarks, providing a simple yet robust framework for step-by-step relational inference.
@@ 6 Future Work
Although ChainMPQ consistently improves the mitigation of relation hallucinations, it still relies on attention distributions as a proxy for visual evidence, which may not fully capture the underlying reasoning process. Future work will incorporate a causality-based attribution mechanism. By detecting and correcting outputs that contradict the image, it prevents errors from cascading. In addition, we will evaluate causal interventions through counterfactual perturbations and uncertainty-aware thresholds to quantify reductions in error propagation on more dataset.
@@ Appendix A Appendix
A.1 LLM usage
Large language models are used solely for refining writing to improve readability, and not for other aspects such as ideas, methods, or experiments.
A.2 Detailed Algorithm
@@ A.3 Dataset Details
R-Bench: A diagnostic benchmark focused on detecting relationship hallucinations in large vision-language models (LVLMs). It comprises image-level and instance-level yes/no questions that test a model’s ability to identify whether specific relations between objects truly exist. The benchmark categorizes hallucinations into three co-occurrence types—relation–relation, subject–relation, and relation–object—to isolate sources of error. R‑Bench samples are sourced primarily from NoCaps, ensuring minimal training overlap. Experiments show that LVLMs frequently over-rely on commonsense priors, misjudging relations in visually ambiguous or adversarial contexts. As a result, R‑Bench is especially effective for fine-grained analysis of relational reasoning errors, offering insights into model biases and areas for improvement.
@@ A.4 Experiment Implementation Details
Because ChainMPQ constructs multiple guiding questions for each original query, the inference time is about four times longer than that of the base model. Nevertheless, the approach remains lightweight and training free. Using two NVIDIA A6000 GPUs, we were able to process nearly 5,000 samples in under three hours, demonstrating that the method can be executed efficiently without requiring large-scale computational resources.
A.5 Detailed Attention Analysis in Case Study
As shown in Figure 6, ChainMPQ progressively guides the model’s reasoning from object localization to relational understanding. The first two sub-questions help the model accurately identify the man and the surfboard, with attention maps showing a gradual focus on the corresponding regions. The subsequent three sub-questions then shift attention toward the interaction between these objects, enabling the model to reason about their relationship. Compared with the direct-answer setting, where attention remains scattered and misaligned with the visual evidence, ChainMPQ produces more concentrated and semantically coherent patterns. In particular, for the question “What is the man standing on”, the attention distribution highlights the surfboard rather than the water, preventing the model from being misled by the phrasing of the original question. This case illustrates how ChainMPQ not only decomposes complex queries into interpretable steps but also reshapes the model’s attention trajectory, thereby supporting more faithful and accurate relational inference.
@@ A.5 Detailed Attention Analysis in Case Study
Figure 6: For each generated question, the attention map during the model’s answer generation process is shown.

### v2

@@ 5 CONCLUSION
We propose ChainMPQ (Multi-Perspective Questions guided Interleaved Text-image Reasoning Chain), a training-free framework that enhances the ability of LVLMs to recognize relationships by utilizing accumulated textual and visual memories. ChainMPQ first strengthens the relevant visual tokens using keywords extracted from the question. It then constructs a set of complementary questions from different perspectives, each closely centered on the subject, the object, and their relationship. These questions are sequentially fed into the model, with both textual and visual memories from earlier rounds propagated to subsequent ones, thereby forming an interleaved chain of image and text. Experimental results demonstrate that ChainMPQ effectively reduces relation hallucinations and improves factuality across multiple LVLMs and benchmarks, providing a simple yet robust framework for step-by-step relational inference.
@@ 6 Future Work
Although ChainMPQ consistently improves the mitigation of relation hallucinations, it still relies on attention distributions as a proxy for visual evidence, which may not fully capture the underlying reasoning process. Future work will incorporate a causality-based attribution mechanism. By detecting and correcting outputs that contradict the image, it prevents errors from cascading.
@@
Acknowledgments
The work is partially supported by the U.S. National Science Foundation (NSF) Grant CRII 2451683, an NVIDIA Academic Grants Program, a U.S. Bank Academic Research Award, the University of California, Merced, and a UC Merced Faculty Research Award. The views and conclusions are those of the authors and do not necessarily reflect the official policy or position of the U.S. Government.
@@ Appendix A Appendix
A.1 Detailed Algorithm
Algorithm 2 describes the complete execution process of ChainMPQ in detail.
@@ A.2 Dataset Details
R-Bench: A diagnostic benchmark focused on detecting relationship hallucinations in large vision-language models (LVLMs). It comprises image-level and instance-level yes/no questions that test a model’s ability to identify whether specific relations between objects truly exist. The benchmark categorizes hallucinations into three co-occurrence types: relation–relation, subject–relation, and relation–object to isolate sources of error. R‑Bench samples are sourced primarily from NoCaps, ensuring minimal training overlap. Experiments show that LVLMs frequently over-rely on commonsense priors, misjudging relations in visually ambiguous or adversarial contexts. As a result, R‑Bench is especially effective for fine-grained analysis of relational reasoning errors, offering insights into model biases and areas for improvement. We clarify that our experiments use the image-level setup of R-Bench. The image-level and instance-level setups share almost the same images and questions, but the instance-level setup also provides bounding box coordinates and focuses more on the relationship between specific object instances. Since most datasets do not provide bounding boxes, and locating the objects involved in a relation is itself an essential ability of LVLMs, we choose the image-level setup for our main experiments to give a fair and clear evaluation of how ChainMPQ improves relational understanding.
A.3 implementation details of multi-perspective question generation
First, we use an NLP tool to extract the subject S, object O, and relation R from the original question. For example, the question “Does the dog chase a disc in the image?” can be decomposed into the subject “dog”, the object “disc”, and the relation “chase”. We then build two questions to locate the subject and the object. In this example, they are “Q1: Where is a dog in the image?” and “Q2: Where is a disc in the image?”
Next, we apply a masking strategy. In each step, one of the three elements is masked, and the other two are used to form a new question. This helps the model focus on the remaining elements and check the original relation from different views. For this example, after masking and reconstruction, we obtain three new questions: “Q3: What is the dog chasing?”, “Q4: What is the disc being chased by?”, and “Q5: What is the relationship between the dog and the disc?” We use GPT-3.5 Turbo to refine each question so that the wording is clear and grammatically correct.
@@ A.4 Experiment Implementation Details
Regarding the comparative experiments with other methods, we reproduced both baseline methods using the latest public code and their official instructions. We only replaced the base models with LLaVA-1.5-7B and InstructBLIP-7B, and we changed the datasets to MMRel and R-Bench to match our setting. All comparisons are made under a fully fair setup. For Constraint-Aware Prompting Wu et al. (2025a), we used the combined-prompt, which is Prompt 10, as it gives the best accuracy. For Detect-then-Calibrate Zheng et al. (2024), we ran their released scripts without any change to the parameters.
A.5 Experiments on general reasoning benchmarks
To explore whether improved relation reasoning also lead to better general reasoning or broader application outcomes. We also evaluated ChainMPQ on two general multimodal benchmarks, MMBench Liu et al. (2024c) and MME Fu et al. (2025). And the results show that ChainMPQ brings consistent gains not only in relation reasoning but also in broader multimodal abilities.
These tasks cover a broad range of multimodal abilities, such as visual logic reasoning, scene understanding, OCR, and object grounding, and are not limited to relation reasoning. Specifically, MMBench defines three capability levels and twenty fine-grained ability dimensions for evaluating different reasoning and perception skills of MLLMs. MME includes tests for both perception and cognition. The perception part covers coarse- and fine-grained skills such as object existence, counting, location, color recognition, poster recognition, celebrity recognition, scene and landmark identification, and artwork recognition. The cognition part includes commonsense reasoning, numerical reasoning, translation, and code reasoning. It contains fourteen sub-tasks in total.
The MMBench results are shown in Table 3. Here Overall is the weighted average over all sub-tasks and is the most important indicator of general ability. RR (Relation Reasoning) refers to relation reasoning. The MME results are shown in Table 4. Here Overall is the average score across all fourteen tasks. Perception measures detection, counting, color, position, and OCR skills, where we do not expect large gains.Cognition measures high-level semantic understanding such as reasoning, description, and complex question answering.
In sum, these results show that improving relation reasoning does lead to better general reasoning. ChainMPQ brings consistent gains on the Overall scores of both MMBench and MME, especially in the Relation Reasoning and Cognition part, confirming that stronger relation understanding supports broader multimodal reasoning abilities.
Table 3: MMBench results (Overall & Relation Reasoning). Higher is better.
[…]
LLaVA-v1.5-7B 66.5 58.8
+ ChainMPQ 67.8 +1.3 61.3 +2.5
InstructBLIP-7B 63.9 52.5
+ ChainMPQ 65.5 +1.6 55.2 +2.7
Qwen2.5-VL-7B 83.2 78.2
+ ChainMPQ 84.7 +1.5 81.8 +3.6
InternVL3-8B 83.6 82.5
+ ChainMPQ 84.2 +0.6 83.9 +1.4
Table 4: MME results (Overall, Perception, Cognition). Higher is better.
[…]
LLaVA-v1.5-7B 1808.4 1506.2 302.1
+ ChainMPQ 1898.3 +89.9 1545.6 +39.4 352.7 +50.6
InstructBLIP-7B 1391.4 1137.1 254.3
+ ChainMPQ 1484.6 +87.2 1176.0 +38.9 308.6 +54.3
Qwen2.5-VL-7B 2312.1 1698.1 613.9
+ ChainMPQ 2350.8 +38.7 1719.2 +21.1 631.5 +17.6
InternVL3-8B 2422.0 1748.4 673.6
+ ChainMPQ 2457.1 +35.1 1771.4 +23.0 685.7 +12.1
A.6 Detailed Case Study
A.6.1 Attention Analysis in Case Study
As shown in Figure 6, ChainMPQ progressively guides the model’s reasoning from object localization to relational understanding. The first two sub-questions help the model correctly identify the man and the surfboard, and the attention maps show a clear shift toward the regions that match each answer. The following three sub-questions gradually move the attention toward the interaction between these objects, which helps the model understand the relationship more clearly.
Compared with the direct-answer setting, where the attention maps are scattered and often point to irrelevant areas, ChainMPQ produces more focused and meaningful patterns. In particular, for the question “What is the man standing on”, the attention highlights the surfboard instead of the water. This prevents the model from being confused by the wording of the original question.
This case shows that ChainMPQ not only breaks down a complex question into simple and clear steps but also adjusts the attention process in a helpful way. By guiding the model step by step, ChainMPQ improves both the reasoning path and the final answer, which leads to more stable and accurate relational predictions.
@@ A.6.1 Attention Analysis in Case Study
Figure 6: For each generated question, the attention map during the model’s answer generation process using ChainMPQ is shown.
A.6.2 Case Study for competitive category in MMRel
We only showed two case studies in the main text because of space limits. This does not mean that our method works better on action or spatial questions than on other types. ChainMPQ is also effective on competitive cases. In Figure 7, we present an example from the competitive domain. This case shows that ChainMPQ can handle competitive questions with the same level of stability as action and spatial questions. It also suggests that our method is not restricted to one type of relation and can work well across different scenarios.
Figure 7: Case study for competitive category